# Reinforcement Learning Dynamics of Network Vaccination and Hysteresis: A Double-Edged Sword for Addressing Vaccine Hesitancy

### v1

@@ Results
Future directions for this work include exploration of different network structures, the possibility of an imperfect vaccine [16], the influence of behavioral interventions, and introducing heterogeneity among agents. Possibly the least realistic modeling assumption is of a $k$-regular network. Scale-free networks are widely recognized as better representative of human populations, and implementing this model on a scale-free network or a real social network could provide insights of the importance of highly connected individuals in the vaccination dilemma. An imperfect vaccine would also increase the realism of our model, adding the potential effect that an agent pays both the cost of vaccination and infection in the same iteration, leading to increased vaccine hesitancy. Behavioral interventions such as vaccine campaigns or the implementation of social distancing are common public health initiatives that warrant further study. Lastly, agent heterogeneity would also be a step towards a more realistic model [43]. In reality, different people are affected in different ways by vaccines and infections, which shapes their decision making in a way we have not accounted for in the present model.
@@
In sum, we investigate reinforcement learning dynamics of vaccination behavior in social networks, where agents learn from their past experiences only with partial information about their neighborhood. We discover the rich dynamical behavior, particularly the bistability of learned vaccination equilibria and hysteresis effect with respect to varying the vaccination cost and the hyperparameters namely the discount rate. These results highlight the double-edged impact of reinforcement learning with implications for mitigating the persistence of vaccine hesitancy.
@@
All computations were done on doob, a 96-core 3.6GHz AMD Epyc system. Code for the project can be found at https://github.com/amcwhorter/vaccine_games.
@@
Investigation of the full parameter space yielded other interesting results in the interplay of discount rate $d$ and temperature $T$.
Figure A1: Hysteresis with respect to discount rate. Change in equilibrium state lags behind the change in the discount factor. Increasing temperature decreases this discrepancy Parameter values to obtain these results: $\alpha=1,\beta=0.4,\gamma=0.5,r=0.2$
Investigation of the full parameter space yielded other interesting results in the interplay of discount rate $d$ and temperature $T$.
As discount factor increases near 1, the vaccination level does not increase, as expected, but remains constant, jumping up right when $d=1$. Going the other way, vaccination level decreases steadily as discount rate decreases. This phenomenon shrinks and disappears as temperature increases.

### v2

@@ Results
Future directions for this work include the exploration of different network structures, the possibility of an imperfect vaccine [18], the influence of behavioral interventions, and the introduction of heterogeneity among agents. Possibly the least realistic modeling assumption is that of a $k$-regular network. Scale-free networks are widely recognized as better representative of human populations, and implementing this model on a scale-free network or a real social network could provide insight into the importance of highly connected individuals in the vaccination dilemma. An imperfect vaccine would also increase the realism of our model, adding the potential effect that an agent pays both the cost of vaccination and infection in the same iteration, leading to increased vaccine hesitancy. Behavioral interventions such as vaccine campaigns or the implementation of social distancing are common public health initiatives that warrant further study. Lastly, agent heterogeneity would also be a step toward a more realistic model [46]. In reality, different people are affected in different ways by vaccines and infections, which shapes their decision making in a way that we have not accounted for in the present model.
@@
In sum, we investigate the reinforcement learning dynamics of vaccination behavior in social networks, where agents learn from their past experiences only with partial information from their neighborhood. We discover rich dynamical behavior, particularly the bistability of learned vaccination equilibria and the hysteresis effect with respect to varying the vaccination cost. These results highlight the double-edged impact of reinforcement learning with implications for mitigating the persistence of vaccine hesitancy.
@@
All computations were performed on doob, a 96-core 3.6GHz AMD Epyc system. Code for the project can be found at https://github.com/amcwhorter/vaccine_games.
@@
Figure A1: Hysteresis with respect to discount rate. Change in equilibrium state lags behind the change in the discount factor. This forms a hysteresis loop with increasing discount rate (blue) and decreasing discount rate (orange). Increasing temperature decreases this discrepancy Parameter values to obtain these results: $\alpha=1,\beta=0.4,\gamma=0.5,r=0.2$.
Investigation of the full parameter space yielded other interesting results in the interplay of discount rate $d$ and temperature $T$. As discount factor increases (blue) near 1, the vaccination level does not increase, as expected, but remains constant, jumping right when $d=1$. Going the other way (orange), vaccination level decreases steadily as discount rate decreases. This phenomenon shrinks and disappears as temperature increases.